If the calculated correlation coefficient is 0,59 what is the percentage of common variance between the variables?
1. 0,3481%
2. 34,81%
3. 59%
4. 1,18%
1
Expert's answer
2020-08-17T18:36:09-0400

solution


Common variance "=R^2"


Where "R =" correlation coefficient



"R^2=0.59^2=0.3481"

answer: 34.81%